# Ctrl-Room: Controllable Text-to-3D Room Meshes Generation with Layout Constraints

### v1

@@ 3.2 Appearance Generation Stage
Given the layout of an indoor scene, we seek to obtain a proper panorama image to represent its appearance. Instead of incrementally generating multi-view images like (Höllein et al., 2023), we generate the entire panorama at once. We utilize ControlNet (Zhang & Agrawala, 2023) to generate a high-fidelity panorama conditioned by the input 3D scene layout. After getting the scene panorama, we recover the depth map by the method in (Shen et al., 2022) to reconstruct a textured mesh through Possion reconstruction (Kazhdan et al., 2006) and MVS-texture (Waechter et al., 2014).
Fine-tune ControlNet. ControlNet is a refined Stable Diffusion (Rombach et al., 2022) model conditioned on an extra 2D input. To condition ControlNet on the scene layout, we convert the bounding box representation into a 2D semantic layout panorama through equirectangular projection. In this way, we get a pair of RGB and semantic layout panoramic images for each scene. However, the pre-trained ControlNet-Segmentation (github, 2023) is designed for perspective images, and cannot be directly applied to panoramas. Thus, we fine-tune it with our pairwise RGB-Semantic layout panoramas on the Structured3D dataset Zheng et al. (2020). As the volume of Structured3D is limited, we apply several techniques during the fine-tuning to augment the training data, including standard left-right flipping, horizontal rotation, and Pano-Stretch (Sun et al., 2019).
Loop-consistent Sampling. A panorama should be loop-consistent. In other words, its left and right should be seamlessly connected. Although the panoramic horizontal rotation in data augmentation may improve the model’s implicit understanding of the expected loop consistency, it lacks explicit constraints and might still produce inconsistent results. Therefore, we propose an explicit loop-consistent sampling mechanism in the denoising process of the latent diffusion model. As shown in Fig. 3(b), we rotate both the input layout panorama and the denoised image for $\gamma$ degree in the sampling process, which applies explicit constraints for the loop consistency during denoising. A concurrent work (Wu et al., 2023) also uses a similar method for panoramic outpainting. More qualitative results in Appendix Figure.2 verify that our simple loop-consistent sampling method achieves good results without introducing additional learnable parameters.
@@ 3.3 Mask-guided Editing
Figure 4: Mask-guided Editing. After editing the scene bounding box, we derive guidance masks from the changes in the semantic layout panoramas. We fill in unoccluded regions and optimize the DIFT (Tang et al., 2023b) features to keep the identity of moved objects unchanged.
@@
A user can modify the generated 3D room by changing the position, semantic class, and size of object bounding boxes. Our method will update the panorama accordingly to reflect the user’s editing in 3D space, which achieves zero-shot editing without expensive re-training.
@@
The editing should achieve two goals, i.e. altering the content according to the user’s input, and maintaining appearance consistency for scene objects. We propose a mask-guided image editing as illustrated Fig. 4, where a chair’s position is moved. In the following, we will explain our method with this example. We denote the semantic panorama from the edited scene as $P_{\rm edited}$, then we derive the guidance masks based on its difference from the original one $P_{\rm ori}$. The source mask $\mathbf{m}_{\rm src}$ shows the position of the original chair, and the target mask $\mathbf{m}_{\rm tar}$ indicates the location of the moved chair, and the inpainting mask $\mathbf{m}_{\rm inpaint}=\{m|m\in\mathbf{m}_{\rm src}\;{\rm and}\;m\notin\mathbf{m}_{\rm tar}\}$ is the unoccluded region. Given these guidance masks, our method includes two steps: the inpainting step and the optimization step. We first fill in the inpaint area by feeding the inpaint mask $\mathbf{m}_{\rm inpaint}$ and edited semantic panorama $P_{\rm edited}$ to the inpainting step. Then, in our optimization step, we optimize the DIFT (Tang et al., 2023b) feature to maintain the visual consistency of relocated objects.
@@
Optimization Step. When the user moves the position of a furniture item, we need to keep its appearance unchanged before and after the movement. The recent work, DIFT (Tang et al., 2023b), finds the learned features from the diffusion network allow for strong semantic correspondence. Thus, we maintain the consistency between the original and moved furniture by requiring their latent features to be consistent. In particular, we extract latent features $F^{l}_{t}$ of the layer $l$ in the denoising U-Net network, at timestep $t$. Then we construct a loss function using the latent features from source area $\mathbf{m}_{\rm src}$ in source panorama $\mathbf{x^{{\rm ori}}_{0}}$ and target area $\mathbf{m}_{\rm tar}$ in inpainted panorama $\hat{\mathbf{x}}^{\rm new}_{0}$.

### v3

@@ 3.2 Appearance Generation Stage
Given the layout of an indoor scene, we seek to obtain a proper panorama image to represent its appearance. Instead of incrementally generating multi-view images like [12], we generate the entire panorama at once. We utilize ControlNet [48] to generate a high-fidelity panorama conditioned by the input 3D scene layout. After getting the scene panorama, we recover the depth map using the method [32] to reconstruct a textured mesh through Possion reconstruction [14] and MVS-texture [40].
Fine-tuning ControlNet. ControlNet is a refined Stable Diffusion [25] model conditioned on an extra 2D input. To condition ControlNet on the scene layout, we convert the bounding box representation into a 2D semantic layout panorama through equirectangular projection. In this way, we get a pair of RGB and semantic layout panoramic images for each scene. However, the pre-trained ControlNet-Segmentation [9] is designed for perspective images, and cannot be directly applied to panoramas. Thus, we fine-tune it with our pairwise RGB-Semantic layout panoramas on the Structured3D [49]. As the volume of Structured3D is limited, we apply several augmentation techniques for the training data, including standard left-right flipping, horizontal rotation, and Pano-Stretch [36].
Loop-consistent Sampling. A panorama should be loop-consistent. In other words, its left and right should be seamlessly connected. Although the panoramic horizontal rotation in data augmentation may improve the model’s implicit understanding of the expected loop consistency, it lacks explicit constraints and might still produce inconsistent results. Therefore, we propose an explicit loop-consistent sampling mechanism in the denoising process of the latent diffusion model. As shown in Fig. 3 (b), we rotate both the input layout panorama and the denoised image by $\gamma$ degree in the sampling process, which applies explicit constraints for the loop consistency during denoising. A concurrent work [44] also uses a similar method for panoramic outpainting. More qualitative results in  Fig. 9 verify that our simple loop-consistent sampling method achieves good results without introducing additional learnable parameters.
@@ 3.3 Mask-guided Editing
Figure 4: Mask-guided Editing. After editing the scene bounding box, we derive guidance masks from the changes in the semantic layout panoramas. We fill in unoccluded regions and optimize the DIFT [38] features to keep the identity of moved objects unchanged.
@@
A user can modify the generated 3D room by changing the position, semantic class, and size of object bounding boxes. Our method will update the panorama accordingly to reflect the user’s edits in 3D space, achieving zero-shot editing without expensive re-training.
@@
The editing should achieve two goals, i.e. altering the content according to the user’s input, and maintaining appearance consistency of the scene objects. We propose a mask-guided image editing as illustrated Fig. 4, where a chair’s position is moved. In the following, we will explain our method with this example. We denote the semantic panorama from the edited scene as $P_{\rm edited}$, then we derive the guidance masks based on its difference from the original one $P_{\rm ori}$. The source mask $\mathbf{m}_{\rm src}$ shows the position of the original chair, and the target mask $\mathbf{m}_{\rm tar}$ indicates the location of the moved chair, and the inpainting mask $\mathbf{m}_{\rm inpaint}=\{m|m\in\mathbf{m}_{\rm src}\;{\rm and}\;m\notin\mathbf{m}_{\rm tar}\}$ is the unoccluded region. Given these guidance masks, our method includes two steps: the inpainting step and the optimization step.
We use $\mathbf{x^{{\rm ori}}_{0}}$ to denote the original image. During the inpainting step, we replace pixels outside the inpainting mask $\mathbf{m}_{\rm inpaint}$ with $\mathbf{x^{{\rm ori}}_{t}}$ and store $\mathbf{m}_{\rm inpaint}$ based on the edited semantic panorama $P_{\rm edited}$. This straightforward approach ensures that the region outside the mask remains unchanged and the area inside the mask is accurately inpainted. We design an optimization step to maintain the appearance of furniture before and after movement and rescaling operations. Recent work DIFT [38], has shown that learned features from the diffusion network enable strong semantic correspondence. Therefore, we ensure consistency between the original and moved furniture by requiring their latent features to be consistent. For more details of the Inpainting Step and Optimization Step, please refer to our supplementary file, Sec.1.2.
4 Experiments
We evaluate Ctrl-Room on three tasks: layout generation, panoramic image generation, and 3D Room generation. We first describe the experimental settings and then validate our method by comparing it with previous methods quantitatively and qualitatively. We further show various scene editing results to demonstrate the flexible control of our method.
4.1 Experiment Setup
Dataset: We train and evaluate our method on the 3D indoor scene dataset Structured3D [49], which consists of 3,500 houses with 21,773 rooms designed by professional artists. Photo-realistic 2D renderings, the labeled room layout, and 3D bounding boxes of furniture are provided in each room. We parse these boxes for typical indoor rooms, such as the bedroom, kitchen, living room, study, and bathroom. Then, we follow [42] to generate text prompts describing the scene layout. Please refer to Appendix Sec.1.3 for more details about data preprocessing. The filtered dataset for training and evaluation consists of 4,961 bedrooms, 1,848 kitchens, 3,039 living rooms, 698 studies, and 1500 bathrooms. For each room type, we use $80\%$ of rooms for training and the remaining for testing. Following DiffuScene [37], we further qualitatively evaluate our layout generation on the 813 living rooms and 900 dining rooms of 3D-FRONT dataset [8].
Figure 5: Qualitative comparison with previous works. For each method, we show a textured 3D mesh in the first row and two rendered images in the second row.
Metrics: Follow previous work [22, 37], Frechet Inception Distance (FID) [11] and Kernel inception distance (KID) [3] are used to measure the plausibility and diversity of 1,000 synthesized scene layouts. We choose FID, CLIP Score (CS) [24], and Inception Score (IS) [28] to measure the image quality of generated panoramas. We also compare the time cost to synthesize an RGB panorama of size $512\times 1024$. To compare the quality of 3D room models, we follow Text2Room [12] to render images of the 3D room model and measure the CLIP Score (CS) and Inception Score (IS). We further conduct a user study and ask $61$ users to score Perceptual Quality (PQ) and 3D Structure Completeness (3DS) of the final room mesh on scores ranging from 1 to 5.
4.2 Comparison with Previous Methods
4.2.1 Qualitative Comparison
Fig. 5 shows some results generated by different methods. The first row shows a textured 3D room model, and the second row shows some perspective renderings from the room model. As we can see, Text2Li-ght [6] cannot generate a reasonable 3D indoor scene. It even fails to ensure the loop consistency of the generated panorama, which leads to distorted geometry and room model. Both MVDiffusion [39] and Text2Room [12] can generate vivid local images as demonstrated by the perspective renderings in the second row. But they fail to capture the more global scale room layout. Similar effects can be seen from the Fig. 1 (a). These two methods often repeat a dominating object, e.g. a bed in the bedroom or fireplace in the living room, multiple times at different places and violate the room layout constraint. In comparison, our method does not suffer from these problems and generates high-quality results. More examples are provided in the Appendix.
Figure 6: Text-conditioned layout generation on Structured3D. Given the text prompt, our method synthesizes a plausible scene layout that matches the description. The generated layout is represented using different colors to indicate various object categories, such as blue for the sofa and brown for the chair.
4.2.2 Layout Generation
Table 1: Quantitative Comparison of layout generation on 3D-FRONT. Note that DiffuScene-w-SC uses an additional network to learn a Shape Code for each furniture, facilitating the evaluation process to retrieve a more accurate CAD model for each furniture. Nevertheless, our method outperforms others in the common settings, where only the generated semantic class and size of each furniture are used for retrieval.
Method Retrieval from Livingroom Diningroom
FID $\downarrow$ KID $\downarrow$ SCA FID $\downarrow$ KID $\downarrow$ SCA
DiffuScene-w-SC [37] Shape Code 35.27 0.64 54.69 32.87 0.57 51.67
ATISS [22] Semantic Bounding Box 40.45 4.57 63.48 36.61 1.90 55.44
DiffuScene-wo-SC [37] Semantic Bounding Box 38.55 1.33 63.54 36.47 1.8 57.04
Ours Semantic Bounding Box 36.0 1.4 56.42 34.78 1.3 54.37
Fig. 6 verifies that our layout generation results are plausible and can offer reliable 3D room layout constraints for the following appearance generation stage. As shown in  Fig. 6, our text-conditioned layout generation module can synthesize natural and diverse typical indoor scenes on the testing set of Structured3D. The size and spatial location of the furniture are reasonable, and the relative positions between the furniture pieces are accurately recovered. Additional objects not described in the text are automatically generated according to the scene prior. We also perform additional experiments using free-style text prompts generated by GPT-4V [46] (Appendix Sec.7.8).
[…]
4.2.3 Panorama Generation
Figure 7: Qualitative comparison for panorama generation. More results are available in the Appendix.
Table 2: Quantitative Comparison of panorama and mesh generation.
Method Panorama Metrics 2D Rendering Metrics 3D Mesh User Study
[…]
Text2Light [6] 56.22 21.45 4.198 81.56 - - 2.732 2.747
MVDiffusion [39] 34.76 23.93 3.21 208.5 - - 3.27 3.437
Text2Room [12] - - - $\geq 9,000$ 25.90 2.90 2.487 2.588
Ours 21.02 22.19 3.56 61.1 25.97 3.14 3.89 3.746
Fig. 7 qualitatively evaluates our generated panoramic images. In Fig. 7, the generated panorama is visualized in a panoramic image viewer to facilitate the user to check the global content. The left side of each column is two zoom-in views, and the right side is the fisheye view. Text2Light [6] suffers from serious inconsistency on the borders of the generated panorama. It also shows a lot of unexpected objects in the image. MVDiffusion [39] suffers from repetitive furniture and fails to synthesize reasonable content for the target room type. In contrast, our method obtains a plausible layout and vivid panorama from the given text prompt.
Table 2 provides quantitative evaluations. We follow MVDiffusion [39] to crop perspective images from the generated panoramas on the test split and evaluate the FID, CS, and IS scores on the cropped multi-view images. In the left part of Table 2, our method achieves the best score in FID, which indicates that our method can better capture the room appearance because of its faithful recovery of the room layout. However, our score on CS is slightly lower than MVDiffusion, which seems insensitive to the number of objects and cannot reflect the quality of room layouts. The IS score depends on the semantic diversity of the cropped images as captured by an image classifier. Text2Light has the best IS score, since its generated indoor scenes often contain unexpected objects.
Fig. 9 shows the performance of our panorama generation module with and without loop-consistent sampling mechanism, indicating the loop-consistent sampling helps the generated panorama obtain better texture consistency.
In terms of running time, our method takes the shortest time. On average, our method takes only 61 seconds to generate a panorama, and another 20 seconds to generate the textured 3D mesh. In comparison, MVDiffusion takes 208 seconds, about 3.5 times longer, to synthesize a panorama. Text2Room needs at least 150 minutes to finish a textured 3D room generation.
4.2.4 3D Room Generation
We then compare the 3D room models in terms of their rendered images. Because of the expensive running time of Text2Room [12], we only test on 12 examples for this comparison. In this comparison, we further skip Text2light and MVDiffusion since we have compared them on panoramas. As the room layout is better captured with a large FOV, we render 60 perspective images of each scene with a $140^{\circ}$ FOV and evaluate their CS and IS scores respectively. Please refer to Appendix Sec.1.4.1 for more details. The results of this comparison are shown in the middle of Table 2. Our method obtains better scores on both metrics than Text2Room.
We further evaluate the quality of the textured 3D mesh model by user studies. For those panorama generation methods, we utilize the depth estimation work [32] to reconstruct a textured 3D room mesh. Further implementation details can be found in the Appendix Sec.1.4. The results of the user study are shown on the right of Table 2. Users prefer our method over others, for its clear room layout structure and furniture arrangement.
Figure 8: Editing examples. (a) resize the TV, (b) replace the chair with a new one.
4.3 Interactive Scene Editing
We demonstrate the scene editing capability of our method in  Fig. 8. In this case, we resize the TV and replace the chair in the generated results. Fig. 1 (b) shows two additional examples of replacing the TV and TV stand and moving the picture upwards. Our method can keep the visual appearance of the moved/resized objects unchanged after editing. More examples can be found in the appendix.
Figure 9: Ablation of loop-consistent sampling examples. We rotate the generated panorama by $180^{\circ}$ to better visualize the leftmost and rightmost content consistency.
5 Conclusion
We present Ctrl-Room, a flexible method to achieve structurally plausible and editable 3D indoor scene generation. It consists of two stages, the layout generation stage and the appearance generation stage. In the layout generation stage, we design a scene code to parameterize the scene layout and learn a text-conditioned diffusion model for text-driven layout generation. In the appearance generation stage, we fine-tune a ControlNet model to generate a vivid panorama image of the room with the guidance of the layout. Finally, a high-quality 3D room with a structurally plausible layout and realistic textures can be generated. We conduct extensive experiments to demonstrate that Ctrl-Room outperforms existing methods for 3D indoor scene generation both qualitatively and quantitatively, and supports interactive 3D scene editing.
6 Limitation
There are still some limitations of Ctrl-Room. Firstly, we only support single-room generation, thus we cannot produce large-scale indoor scenes with multiple rooms. Secondly, the generated 3D room still contains incomplete structures in invisible areas because of the occlusion and poor performance of the panoramic depth estimator. A promising direction is to learn a text-driven diffusion model to produce one or more RGB-D panorama images under the scene layout constraints. Lastly, as we explore injecting 3D scene information into pretrained 2D models, thus we rely on 3D labeled scene dataset to drive the learning and fine-tuning process. Leveraging scene datasets with only 2D labels to learn 3D priors is also a promising direction. We leave the aforementioned limitations for our future efforts.
7 Supplymentary File
In the supplementary file, we first present more details about our scene code diffusion model in Sec. 8, then we elaborate the mask-guided editing method in Sec. 9. Next, we provide our dataset pre-processing, text prompt generation, and implementation details in Sec. 10 and Sec. 11 respectively. Additional experiment results are also illustrated, including panorama generation comparisons in Sec. 12, room mesh comparisons in Sec. 13 and user studies in Sec. 14. Furthermore, we demonstrate that our scene code diffusion model can be trained with free-style text prompts in Sec. 15.
8 Scene Code Denoising Network
In the Layout Generation Stage, we use a holistic scene code to parameterize the indoor scene and design a diffusion model to learn its distribution. Specifically, given a 3D scene $\mathcal{S}$ with $N$ objects, we represent the scene layout as a holistic scene code $\mathbf{x_{0}}=\{\mathbf{o_{i}}\}^{N}_{i=1}$. We encode each object $o_{i}$ as a node with various attributes, i.e., center location $l_{i}\in\mathbb{R}^{3}$, size $s_{i}\in\mathbb{R}^{3}$, orientation $r_{i}\in\mathbb{R}$, class label $c_{i}\in\mathbb{R}^{C}$. Each node is characterized by the concatenation of these attributes as $\mathbf{o}_{i}=[c_{i},l_{i},s_{i},r_{i}]$. As shown in Fig. 10, our scene code denoising network of the layout diffusion model is built upon IDDPM [21]. The whole architecture of the layout diffusion model is similar to IDDPM, while we replace the upsample and downsample blocks with 1D-convolution network in the U-Net, and insert attention blocks after each residual block to capture both the global context among objects and the semantic context from the input text prompt. The input encoding head processes different encoding of the node attributes, e.g., semantic class labels, box centroid, and box orientation. After adding noise, the input encoding is fed into the U-Net to obtain a denoised scene code. During the forward phase, as in IDDPM, we iteratively perform the denoising process and generate a scene code from a partial scene textual description.
We further investigate how the physical regularization term impacts the final 3D scene layout. In Fig. 11, we use two text prompts for layout generation our layout diffusion model trained with and without our physical regularization term, respectively. As can be seen, the diffusion model trained with the physical violation loss can effectively reduce the occurrence of furniture penetrating walls, and also help to regulate the orientation of the sampled furniture, resulting in more reasonable layouts than the model without the physical regularization term.
Figure 10: The detailed structure of the scene code denoising network. We here take the bedroom for example to demonstrate the dataflow of the scene code denoiser. The scene code tensor $\mathbf{x_{0}}\in\mathbb{R}^{N\times D}$, where $N=23,D=32$.
Figure 11: Ablation study of the physical violation loss. Two text prompts of study are used for layout generation using our diffusion model trained without $\mathcal{L_{\rm physical}}$ (left) and with $\mathcal{L_{\rm physical}}$ (right), respectively. As a result, in the left sample, diffusion model without $\mathcal{L_{\rm physical}}$ generates a green desk that penetrates the wall. In the right sample, this phenomenon is alleviated and regulated after using the physical violation loss. Note that the sampling results of these two versions of diffusion models are slightly different since the denoise distribution is different even given the same text prompt.
9 Mask-Guided Editing
To achieve consistent and seamless 3D scene editing, it should achieve two goals, i.e. altering the content according to the user’s input, and maintaining appearance consistency for scene objects. We propose a mask-guided image editing as illustrated Fig.4 in the main paper, where a chair’s position is moved. In the following, we will explain our method with this example. We denote the semantic panorama from the edited scene as $P_{\rm edited}$, then we derive the guidance masks based on its difference from the original one $P_{\rm ori}$. The source mask $\mathbf{m}_{\rm src}$ shows the position of the original chair, and the target mask $\mathbf{m}_{\rm tar}$ indicates the location of the moved chair, and the inpainting mask $\mathbf{m}_{\rm inpaint}=\{m|m\in\mathbf{m}_{\rm src}\;{\rm and}\;m\notin\mathbf{m}_{\rm tar}\}$ is the unoccluded region. Given these guidance masks, our method includes two steps: the inpainting step and the optimization step. We first fill in the inpaint area by feeding the inpaint mask $\mathbf{m}_{\rm inpaint}$ and edited semantic panorama $P_{\rm edited}$ to the inpainting step. Then, in our optimization step, we optimize the DIFT [38] feature to maintain the visual consistency of relocated objects.
@@ 9 Mask-Guided Editing
Optimization Step. When the user moves the position of a furniture item, we need to keep its appearance unchanged before and after the movement. The recent work, DIFT [38], finds the learned features from the diffusion network allow for strong semantic correspondence. Thus, we maintain the consistency between the original and moved furniture by requiring their latent features to be consistent. In particular, we extract latent features $F^{l}_{t}$ of the layer $l$ in the denoising U-Net network, at timestep $t$. Then we construct a loss function using the latent features from source area $\mathbf{m}_{\rm src}$ in source panorama $\mathbf{x^{{\rm ori}}_{0}}$ and target area $\mathbf{m}_{\rm tar}$ in inpainted panorama $\hat{\mathbf{x}}^{\rm new}_{0}$.